# Dark Matter from Holography

### v1

@@ Abstract
The model we have presented here provides a novel mechanism to generate dark matter, and it leads to the generation of a positive effective cosmological constant from a preexisting negative $\rho_{\Lambda}$. Further, it provides a natural mechanism for the rough equivalence of the baryon and dark matter densities. The key unresolved issue is whether such a form for dark matter can cluster as observed. A resolution of this question would require the development of an underlying Lagrangian for the holographic model.
@@
$\mathcal{N}(L)\sim\exp\left[\frac{A(L)}{4}\right]$ (13)
@@
$\rho_{\rm holo}(L)\propto L^{-2}$ (14)
which is the same structural dependence that underlies the HDM construction in Eq. (3). Thus, HDM can be viewed physically as the bulk gravitational response to the finite number of microstates allowed by the cosmic horizon, rather than a cloud of microscopic particles, aligning its physical origin squarely with quantum-gravitational microstructure.
@@
$S_{\rm matter}(L)+S_{\rm grav}(L)\leq\frac{A(L)}{4}$ (15)
which does not privilege dark energy over dark matter in any way. Therefore, assigning part of the holographically allowed energy budget to a clustering sector is entirely consistent with these QG principles. Once this horizon-governed energy is projected into the bulk Einstein equations, it naturally behaves as a matter component with scaling determined by the infrared cutoff precisely as realized in the Ricci HDM model we have presented.

### v2

@@ 5 Conclusions and Discussion
The model we have presented here provides a novel mechanism to generate dark matter, and it leads to the generation of a positive effective cosmological constant from a preexisting negative $\rho_{\Lambda}$. Further, it provides a natural mechanism for the rough equivalence of the baryon and dark matter densities. The key unresolved issue is whether such a form for dark matter can cluster as observed. While we have provided a reasonableness argument in the previous section, a complete resolution of this question would require the development of an underlying Lagrangian for the holographic model.
@@
$\mathcal{N}(L)\sim\exp\left[\frac{A(L)}{4}\right]$ (5.1)
@@
$\rho_{\rm holo}(L)\propto L^{-2}$ (5.2)
which is the same structural dependence that underlies the HDM construction in Eq. (2.1). Thus, HDM can be viewed physically as the bulk gravitational response to the finite number of microstates allowed by the cosmic horizon, rather than a cloud of microscopic particles, aligning its physical origin squarely with quantum-gravitational microstructure.
@@
$S_{\rm matter}(L)+S_{\rm grav}(L)\leq\frac{A(L)}{4}$ (5.3)
which does not privilege dark energy over dark matter in any way. Therefore, assigning part of the holographically allowed energy budget to a clustering sector is entirely consistent with these principles. Once this horizon-governed energy is projected into the bulk Einstein equations, it naturally behaves as a matter component with scaling determined by the infrared cutoff precisely as realized in the Ricci HDM model we have presented.